# Dolphin: A Programmable Framework for Scalable Neurosymbolic Learning

### v1

@@ 4.4 RQ2: Accuracy
We observe that in all cases, Dolphin achieves state-of-the-art accuracy, except in CLUTRR, where Dolphin’s accuracy is slightly lower than that of Scallop. For most Medium and Large versions of the benchmarks, Dolphin achieves better accuracy, whereas the baselines either report lower accuracy due to the complexity of the benchmark (e.g., HWF) or fail to converge within the time limit (e.g., Scallop on PathFinder-Large). Most importantly, these results show that Dolphin’s scalability improvements do not come at the cost of accuracy.
@@ 4.5 RQ3: Provenance Comparisons
We perform ablation studies to compare the effectiveness of the DAMP and DTKP-AM provenances for each benchmark. Figure 5 presents the accuracy of Dolphin using DAMP and DTKP-AM in the top graph, and the average training time per epoch in the bottom. We observe that the relative effectiveness of the provenances differ across benchmarks. For instance, in the MNIST-SumN benchmark, the DAMP provenance is more effective than the DTKP-AM provenance, since the top-k proofs are not able to capture all the possible ways in which the sum of the digits can be computed. In contrast, for HWF, the DTKP-AM provenance is more effective than DAMP since the tags in DAMP provenance lack the structure needed to capture the semantics of the symbolic program. Otherwise, DTKP-AM mostly has a slight edge over DAMP. On the other hand, DAMP is faster than DTKP-AM for all benchmarks, since it does not track individual proofs. We thus see that selecting provenances is a trade-off between accuracy and training time.
@@ 5 Related Work
Neurosymbolic programming frameworks. Frameworks like Scallop (Li et al., 2023), DeepProbLog (Manhaeve et al., 2018), and ISED (Solko-Breslin et al., 2024) provide a simple interface for neurosymbolic programming. There are also domain-specific tools like NeurASP Yang et al. (2023) for answer set programming and NeuralLog Chen et al. (2021) for phrase alignment in NLP. While these frameworks provide intuitive abstractions, they are bottle-necked due to expensive data transfers between symbolic computations done on CPU versus neural computations that execute on GPU, making neurosymbolic learning hard to scale. In contrast, Dolphin provides a deeper integration of the two worlds by building a Python-based API on top of PyTorch, which scales better.
Scalability improving techniques. Several optimization techniques have been proposed to improve the scalability of differentiable reasoning algorithms. For instance, Greedy NTP (Minervini et al., 2020a) reduces the computation cost of Neural Theorem Prover (NTP) (Rocktäschel & Riedel, 2017) by tracking only a subset of proof states using nearest neighbor search. Likewise, the conditional theorem prover (Minervini et al., 2020b) employs a machine learning-based proof selection technique. However, unlike Dolphin, these methods are point solutions that do not fundamentally address the scalability challenge for neurosymbolic learning.
Specialized neurosymbolic solutions. There are many specialized solutions for various neurosymbolic tasks. For instance, NGS (Li et al., 2020) uses a hand-coded syntax to specify the structure of mathematical expressions for HWF. More general solutions, such as NS-CL (Mao et al., 2019) includes a framework for visual question answering that learns symbolic representations for text and images. NeRd (Chen et al., 2021) transforms questions in natural language into executable programs based on symbolic information extracted from text. In this work, we rather focus on building a general programming framework that improves the scalability of diverse neurosymbolic programs.
@@ 6 Conclusion and Limitations
We proposed Dolphin, a programmable framework for scaling neurosymbolic learning. Dolphin provides abstractions for writing symbolic programs along with pluggable vectorized provenances to compute symbolic gradients. This allows users to write differentiable symbolic programs in Python within PyTorch pipelines that can scale to complex programs and large datasets. We show that Dolphin scales significantly better than existing neurosymbolic frameworks while achieving state-of-the-art performance on a variety of tasks.
@@
A limitation of Dolphin is that it requires the user to express symbolic programs in a batched manner. While this is a common pattern within deep learning, it may be restrictive to users new to batched programming. Another limitation is that while Dolphin works well with most models, the representation of Distributions needed by generative models (e.g., Causal LLMs like Llama) has not yet been investigated. We leave this to future work.
Appendix A Dolphin Experiment Details for Benchmarks
@@
The following are more details on the experiment setup. For each experimental trial, we report the highest evaluation accuracy over all epochs. Unless otherwise noted, each trial was run on a machine with Intel Xeon Gold 6248 (2.50 GHz) CPUs and NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs.
A.1 MNIST Sum-N
@@
Each of the MNIST Sum-N tasks had a batch train size of 64 samples, a learning rate of 0.001, and a top-k value of 1. Each of the tasks were trained on a dataset size of the original MNIST dataset divided by N of Sum-N. Sum5’s dataset consisted of 12000 train samples and 2000 test samples. Sum10’s dataset consisted of 6000 train samples and 1000 test samples. Sum15’s dataset consisted of 4000 train samples and 666 test samples.
A.2 Hand-Written Formula
@@
For Hand-Written Formula we trained each task with a batch train size of 64 samples. The learning rate was 0.0001, sample-k was 7, and top-k value was 3. Length 7’s dataset consisted of 9600 samples for training, 2400 samples for testing. Length 15 consisted of 24000 training samples and 6000 testing samples. Length 19 consisted of 32000 training samples and 8000 testing samples.
A.3 PathFinder
@@
For each of the PathFinder tasks, we used a batch train size of 64 samples, a learning rate of 0.0001, and a top-k value of 3. Each task’s dataset consisted of 539459 images for training and 59940 images for testing. Each task had its own dataset of images with dimensions of the task’s pixel size.
A.4 CLUTRR

### v2

@@ 3.4 RQ2: Accuracy
We observe that in all cases, Dolphin achieves state-of-the-art accuracy among general-purpose neurosymbolic frameworks, except in CLUTRR-3, where Dolphin’s accuracy is slightly lower than Scallop’s. As we scale up to larger versions of the benchmarks, Dolphin achieves better accuracy, because the baselines either report lower accuracy due to the complexity of the benchmark (e.g., black-box sampling techniques such as ISED/IndeCateR+ on HWF) or fail to converge within 10 hours due to slower per-epoch train time (e.g., Scallop on PathFinder-256). As a result, Dolphin significantly outperforms the second-best benchmark on the largest dataset versions, achieving a 20% gain on HWF, 33% on Path, and up to 38% on Mugen. Most importantly, these results show that not only do Dolphin’s scalability improvements not come at the cost of accuracy, but Dolphin enables SOTA accuracy when previously unattainable.
@@ 3.5 RQ3: Provenance Comparisons
We perform ablation studies to compare the effectiveness of the DAMP and DTKP-AM provenances for each benchmark. We share the graphs in Figure 8 (Appendix E). In all cases, training with the DAMP provenance takes around 24.19 seconds per epoch less than with DTKP-AM on average.
However, the effectiveness of each provenance varies across benchmarks. For all variations of Path, CLUTRR, and Mugen, both provenances achieve comparable accuracies, with DTKP-AM having a slight edge. For SumN, DAMP provenance is more effective than the DTKP-AM by 72% points on average, since the top-k proofs cannot capture all the possible ways in which sums of digits can be computed.
In contrast, for HWF, DTKP-AM is more effective than DAMP by an average of 42.2% points. Each step of the HWF program, shown in Appendix H, involves both a concatenation operation and a partial parsing operation before the final expression is evaluated to produce a result. As such, it is difficult for the tags in DAMP to capture the semantics of the symbolic program. In the case of DTKP-AM, each tag is a collection of proofs over input symbols corresponding to logits derived from the neural model. Therefore, any calculated gradients can be directly backpropagated to the logits that most influenced the output, making this a more effective provenance for this task.
@@ 4 Related Work
Neurosymbolic programming frameworks. Apart from existing frameworks like Scallop (Li et al., 2023), DeepProbLog (Manhaeve et al., 2018), and ISED (Solko-Breslin et al., 2024), there exist domain-specific tools such as NeurASP (Yang et al., 2021) for phrase alignment. These approaches often suffer from inefficiencies due to CPU-bound symbolic reasoning.
Scaling techniques. Various methods exist to scale differentiable reasoning. LYRICS (Marra et al., 2019), Logic Tensor Networks (Badreddine et al., 2022), and Tensorlog (Cohen et al., 2020) compile first-order logic constraints into GPU-compatible computation graphs. Other techniques, such as Greedy NTP (Minervini et al., 2020a) and the conditional theorem prover (Minervini et al., 2020b), optimize proof search using heuristics. However, these methods are often task-specific and lack generalizability to broader neurosymbolic learning, or they struggle to scale due to memory constraints when grounding symbolic computations on GPU.
Specialized neurosymbolic solutions. There are many specialized solutions for various neurosymbolic tasks. For instance, NGS (Li et al., 2020) uses a hand-coded syntax to specify the structure of mathematical expressions for HWF. More general solutions, such as NS-CL (Mao et al., 2019) includes a framework for visual question answering that learns symbolic representations for text and images. NeRd (Chen et al., 2021) transforms questions in natural language into executable programs based on symbolic information extracted from text. (Orvieto et al., 2023) proposes a recurrent neural network architecture that achieves 95% accuracy on Path (S) and 94% on Path (M). In contrast, Dolphin is a general system that tries to scale diverse neurosymbolic programs.
@@ 5 Conclusion and Limitations
We proposed Dolphin, a system for scaling neurosymbolic learning. Dolphin provides abstractions for writing symbolic programs along with pluggable vectorized provenances to compute symbolic gradients. This allows users to write differentiable symbolic programs in Python within PyTorch pipelines that can scale to complex programs and large datasets. We show that Dolphin scales significantly better than existing neurosymbolic frameworks while achieving state-of-the-art performance on a variety of tasks.
@@
A limitation of Dolphin is that it needs the user to write programs in a batched manner. This is a common pattern within deep learning but may be restrictive to users new to batched programming. Also, while Dolphin works well with most models, the representation needed by generative models (e.g., Causal LLMs) has not been investigated yet. Dolphin also lacks support for non-deterministic symbolic programs. We leave these for future work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A DTKP-AM Provenance
We clarify and expand on some aspects of the DTKP-AM provenance.
A.1 WMC approximation
In this section, we emphasize that DTKP-AM does not perform precise weighted model counting (WMC) and address possible shortcomings that could arise. A hardware-efficient vectorization of exact WMC is beyond the scope of this paper, and is itself an active area of research. Instead, we use the following add-mult approximation of WMC:
$\Pr(t)=\sum_{i}\Pr(t_{i})=\sum_{i}\prod_{j}\text{norm}(t_{ij})$
We note that this approximation upper bounds the result from DTKP-WMC: the coarseness arises from the summation, which may double count models that satisfy more than one of the proofs. However, add-mult achieves significant computational speedup since it simplifies the exponential enumeration over all possible models into a linear pass over the tag’s elements.
We further claim that this approximation does not destroy all the semantics from DTKP-WMC due to DTKP-AM’s faithful implementation of the semiring operations $\oplus$ and $\otimes$ for tracking top-k proofs. DTKP-AM tags therefore remain similar to DTKP-WMC tags at every intermediate symbolic reasoning step. By contrast, the imprecise add-mult is a one-time transformation of the final tags into probabilities, performed only after the tags have been propagated through the entire symbolic program. Crucially, we show there exists information that is uniquely captured by top-k tag operations, and is not lost when fuzzily converting the tags to probabilities.
As a simple illustrative example, consider using Apply with the following toy function:
$f(a,b)=\begin{cases}\mathbb{T}&a=b\\ \mathbb{F}&\text{otherwise}\end{cases}$
For any distribution $D$ of mutually exclusive input symbols (e.g. the digit classification of a CNN), we intuitively would like the distribution $f(D,D)$ to assign a probability of 1 to symbol $\mathbb{T}$ and a probability of 0 to symbol $\mathbb{F}$. According to our semantics, the tag for $\mathbb{T}$ is actually given by:
$f(D,D)(\mathbb{T})=\bigoplus_{i}(D(i)\otimes D(i))$
However, if we were to use DAMP to compute the tags for $f(D,D)$, the provenance treats the two input distributions as independent when they are the exact same distribution! Thus, the probability assigned to $\mathbb{T}$ by $f(D,D)$ is incorrectly calculated as:
$f(D,D)(\mathbb{T})=\sum_{i}(D(i))^{2}$
On the other hand, consider any top-k provenance that satisfies:
$t\otimes t^{\prime}=\text{top}_{k}(\{t_{i}\cup t^{\prime}_{j}\mid(t_{i},t^{\prime}_{j})\in t\times t^{\prime}\})$
where $\times$ is the set Cartesian product. Note that DTKP-AM does satisfy this condition, where the set union is implemented with an element-wise minimum. Now assuming $D(i)$ is initialized in the natural way (i.e. a tag consisting of a single proof containing just the input symbol $i$), then $D(i)\otimes D(i)=D(i)$ and therefore:
$f(D,D)(\mathbb{T})=\bigoplus_{i}D(i)$
Under both and the add-mult approximation, the probability of $\mathbb{T}$ is:
$\sum_{i}\prod_{j}\text{norm}(D(i)_{ij})=\sum_{i}D(i)_{ii}=\sum_{i}\Pr(i)=1$
for any normalized $D$ with at most $k$ symbols. Even if the number of symbols exceeds $k$, we note that the distributions we seek to learn are often skewed (an accurate model should assign a probability to the ground truth that significantly outweighs the other symbols). For such distributions, DTKP and DTKP-AM would still yield the same probability for $\mathbb{T}$, and it is much closer to 1 than the sum of squares result from DAMP.
While this example may seem contrived, it still suggests the smaller role a “correct” WMC can have on the final answer compared to $\oplus$ and $\otimes$ implemented with proper set-based semantics. We even hypothesize that in most cases, the add-mult approximation does not meaningfully affect the final result compared to DTKP-WMC. This is empirically demonstrated by our benchmark results, which shows DTKP-AM achieving similar accuracy to Scallop’s implementation of DTKP-WMC. In fact, DAMP can be considered as a sort of ablation, where both the WMC and semiring operations use fuzzy add-mult semantics instead of a set-based one, and indeed, its accuracy often performs worse than both DTKP-WMC and DTKP-AM.
A.2 Role of +$\infty$+ ∞ and $\infty$- ∞
In this section, we motivate the use of +$\infty$+ ∞ and $\infty$- ∞ in DTKP-AM’s tensor representation of tags. Because tensors $t$ are rectangular where every proof $i$ and symbol $j$ must have an entry $t_{ij}$, we require a way to denote the absence of an input symbol from a proof, and the absence of a proof from a tag. Importantly, an absent symbol should not influence the probability of a proof (i.e. its normalized value should contribute 1 to the probability’s product), and an absent proof should not influence the probability of a tag (i.e. it should contribute 0 to the sum during add-mult WMC). Indeed this is captured by our definition of norm, which clamps +$\infty$+ ∞ to 1 (representing absent symbols) and $\infty$- ∞ to 0 (representing absent proofs) during any probability calculation. While this introduces clamping operations, PyTorch’s implementation of clamp backpropagation ensures a gradient of 1 everywhere, even on the clamp boundaries (source: https://github.com/pytorch/pytorch/pull/7049).
Since $\mathbf{\hat{0}}$ corresponds to the tag consisting of no proofs (i.e. a tag with probability 0), we initialize it to be a tensor where every proof is absent (all $\infty$- ∞). Likewise, since $\mathbf{\hat{1}}$ corresponds to the tag consisting of a single empty proof (i.e. a tag with probability 1), we initialize it to be a tensor where every symbol is absent from the first row / proof (all +$\infty$+ ∞), while the remaining rows / proofs are absent (all $\infty$- ∞).
A.3 Further reading
For a more in-depth explanation of provenances in general, including the formalization of DTKP semantics with Boolean formulae, see Section 4 of (Li et al., 2023). For worked examples of provenance computation with comparisons of top-k provenances to DAMP, we refer the reader to (Scallop Language Group, 2022).
Appendix B Control Flows and Recursion in Dolphin
Config Time for UDF (s) Time for Tag Computations (s) Total Time (s)
No Parallelism 36.24 (C) 461.02 (C) 497.26
Parallelized Tag Computations 14.13 (C) 75.125 (G) 89.25
Table 4: Time taken by the symbolic program for the HWF task split by the time spent on the CPU and GPU. UDFs refer to user-defined functions where control flows reside for HWF. The times annotated with C and G indicate time spent on the CPU and GPU, respectively.
1Coord = namedtuple(’Coord’, [’x’, ’y’])
2
3def compute_paths(paths, edges):
4 new_paths = apply_if(paths, edges, lambda p1, p2: Coord(p1.x, p2.y), lambda p1, p2: p1.y == p2.x)
5 merged = union(paths, new_paths)
6 # checking for convergence via fix-point
7 if merged.symbols == paths.symbols:
8 return merged
9 else:
10 return compute_paths(merged, edges)
11
12edges = Distribution(model(img), points)
13paths = compute_paths(edges, edges)
Figure 4: Example of a transitive closure computation in Dolphin.
In this section, we provide a more detailed explanation of how Dolphin handles control flows and recursion. In Dolphin, control flows largely exist within the lambda functions supplied to the ‘Apply‘, ‘ApplyIf‘, and ‘Filter‘ operations, which can be arbitrary Python functions over the symbols in the Distributions. As discussed in Section 2.1.2, these functions can include complex operations like if-then-else branches, loops, and even recursions. We do assume that divergent control flows are resolved within the lambda function itself. The nature of these functions means that they cannot be parallelized over the GPU. Instead, they are executed sequentially on the CPU, while the associated tags are computed parallely on the GPU. We optimize the design of the Distribution class so that there is one set of CPU-based computations for the entire batch of samples rather than one set of computations for each sample, which is typical of other neurosymbolic frameworks. This allows Dolphin to maintain the benefits of parallelism even while the user-defined functions are executed sequentially.
B.1 Control Flow in HWF
We demonstrate this by showing the time taken by the symbolic program for the HWF task split by the time spent on the CPU and GPU in Table 4. The first row shows the time taken when the neurosymbolic model is run sequentially on the CPU with no parallelism. The second row shows the time taken when tag computations are parallelized on the GPU over batches of 64 samples each. The times annotated with C and G indicate time spent on the CPU and GPU, respectively. We only show the time taken in the forward pass in the table.
Observe that the time, both for UDF computation and for Tag computation, decreases as we move from sequential CPU evaluation to the batched evaluation. Due to Dolphin’s design, increases in batch size result in fewer CPU operations, since the set of CPU operations is shared for the entire batch, while parallelizing more tag computations over the entire batch.
B.2 Recursion
In order to write recursive computations in Dolphin, one has two choices: either supply a recursive user-defined function to the Dolphin primitives, or write a more fine-grained program in Python that uses Dolphin primitives in the base case as well as the recursive case, set to terminate once a condition is met. Here, the diverging control flows can be merged using the Union primitive. We follow the latter approach for tasks involving recursion, such as Path and CLUTRR. The crux of those programs involves performing a transitive closure computation over a graph, represented by a set of edges for Path or relations for CLUTRR. We show an example of a transitive closure computation in Figure 4.
Here, lets say that model is a neural model that predicts the edges between each pair of points in a graph, represented by points. The compute_paths function computes the transitive closure of the graph by iteratively applying the edges to the paths. The ApplyIf function applies the edges to the paths if the end of the first path is the same as the start of the second path. The Union function merges the new paths with the existing paths. The function compute_paths is called recursively until a fixpoint is reached, specifically until no new paths can be added. This is a simple example of a recursive computation in Dolphin, and also forms the core program needed for the PathFinder task. We perform a similar recursive computation for the CLUTRR task, where we find the transitive closure of a graph representing relations between people in a passage.
Appendix C On the Language and Semantics
C.1 Language
To develop the operations provided by Dolphin, we studied several neurosymbolic tasks to determine the most common operations needed for these tasks. We found that the main operation needed in most programs is to apply a function to symbols from different input models and relations. This is primarily achieved via the join operation in Datalog, but we introduce the Apply or ApplyIf primitives for a more Pythonic approach. Filters are used to remove symbols violating conditions, similar to Datalog selections, while Union mimics the disjunction operation in Datalog, typically needed for writing recursive programs as described in Appendix B.
C.2 Semantics
We designed Dolphin to be a general-purpose neurosymbolic framework able to support various semantics, as long as they can be expressed as operations over tags tracked via the Distribution class. Dolphin assumes that the provenance supplied to it offers both the conjunction and disjunction operations that operate over combinations of tags from input symbols, as well as a way to translate tags to probabilities. As long as these assumptions are satisfied, the primitives of Dolphin preserve the semantics offered by the provenances.
As such, supplying the DAMP provenance to the Dolphin program introduces basic fuzzy semantics which are preserved by Dolphin. However, there are cases where the independence assumptions may not hold and fuzzy semantics may not be appropriate.
The DTKP-AM provenance, on the other hand, offers an alternative without the assumption of variable independence, except on the input variables. At each step of the program, each symbol is associated with the tags of the input symbols that produce it via the proofs. Again, since DTKP-AM satisfies the aforementioned assumptions, the top-k semantics of the provenance are preserved.
These tags are then translated into probabilities by performing an add-mult operation over the proofs. This approximation of the WMC operation is more complex and results in a more precise translation of tags to probabilities. However, as we see in the experiments where Scallop uses DTKP-WMC, the accuracies achieved by DTKP-AM and DTKP-WMC are comparable.
Appendix D Dolphin Experiment Details for Benchmarks
D.1 Comparison of Per Epoch Training Times
Task Dolphin Scallop LTN ISED IndeCateR+
[…]
SumN-5 10.77 184.76 3.43 4.63 0.09 59.54 59.93 59.54 1.11
SumN-10 10.49 341.57 32.56 OOM 216.54 20.64 32.14 3.06
SumN-15 10.47 493.87 47.17 OOM 653.39 62.41 23.84 2.28
HWF-7 152.87 499.57 3.27 N.A. 790.04 5.17 540.26 3.53
HWF-15 858.8 1.49e4 17.35 6.84e3 7.97 2.51e3 2.93
HWF-19 1.4e3 3.92e5 280 1.05e4 7.5 4.18e3 2.99
Path-32 1.29e3 2.2e3 1.71 N.A.
Path-128 1.67e3 4.18e3 2.5
Path-256 1.97e3 1.13e4 5.74
CLUTRR-3 152.21 429.97 2.82 N.A.
CLUTRR-4 165.13 783.11 4.74
Mugen-1K 165.74 133.68 0.81 N.A.
Mugen-5K 826.31 634.86 0.77
@@ D.1 Comparison of Per Epoch Training Times
The following are more details on the experiment setup. For each experimental trial, we report the highest evaluation accuracy over all epochs. Unless otherwise noted, each trial was run on a machine with Intel Xeon Gold 6248 (2.50 GHz) CPUs and NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs. Additionally, for Scallop, the accuracy of the best-performing provenance is reported. Table 5 shows the training time per epoch for all of the baselines in each of the benchmarks.
Dolphin trains slightly slower than Scallop on both versions of the Mugen benchmark. This is because the Dolphin program written for Mugen uses Python objects and operations that are not fully batchable across samples. In contrast, the Scallop program, which is written in a compiled and optimized language, runs around 1.3x faster than Dolphin on average per iteration. However, Dolphin requires only 20 epochs to converge whereas Scallop requires almost 1000 epochs  (Li et al., 2023). The accuracy curves are shown and discussed in 6 and 7. As a result, Dolphin’s total training time is still lower than Scallop’s (3.5x for Mugen-1K).
D.2 MNIST Sum-N
For this task, the base neural network model is a standard CNN classifying each image into 10 classes of digits (0, 1, $\ldots$, 9). The symbolic module sums the Distribution objects over the logits output by the neural model for each image.
@@ D.2 MNIST Sum-N
Each of the MNIST Sum-N tasks had a batch train size of 64 samples, a learning rate of 0.001, and a top-k value of 1. Each of the tasks were trained on a dataset size of the original MNIST dataset divided by N of Sum-N. Sum5’s dataset consisted of 12000 train samples and 2000 test samples. Sum10’s dataset consisted of 6000 train samples and 1000 test samples. Sum15’s dataset consisted of 4000 train samples and 666 test samples. Figure 7 is a high-level overview of the Sum-N model’s architecture.
Figure 5: Components of the SumNNet model written in Figure LABEL:fig:mnist_sum_n_code.
D.3 Hand-Written Formula
For Hand-Written Formula, the perception model is again a standard CNN that classifies images into 14 classes: 10 digits (0, 1, $\ldots$, 9), and 4 operations (+, -, $\times$, and /). The Dolphin program for this task builds strings of formulae from the outputs of the neural model and evaluates them using Python’s eval function, demonstrating the ability of Dolphin to support black-box functions.
@@ D.3 Hand-Written Formula
We trained each task with a batch train size of 64 samples. The learning rate was 0.0001, the global sampling value was 7, and top-k value was 3. Length 7’s dataset consisted of 9600 samples for training, 2400 samples for testing. Length 15 consisted of 24000 training samples and 6000 testing samples. Length 19 consisted of 32000 training samples and 8000 testing samples.
D.4 PathFinder
For this task, the perception model is also a CNN, but it predicts edges between pairs of nodes (denoted by dashes) as well as the end points depicted in the image of the maze. The Dolphin program for this task is recursive since it must search for paths between the two dots.
@@ D.4 PathFinder
For each of the PathFinder tasks, we used a batch train size of 64 samples, a learning rate of 0.0001, and a top-k value of 1. Each task’s dataset consisted of 539459 images for training and 59940 images for testing. Each task had its own dataset of images with dimensions of the task’s pixel size.
D.5 CLUTRR